EXPERIMENTAL DETERMINATION OF THE BINDING ENERGIES OF THE T-SHAPED AND LINEAR ISOMERS OF THE Ne⋯ICl(X,v′′=0) VAN DER WAALS COMPLEX

\underline{DAVID B. STRASFELD}, JOSHUA P. DARR and RICHARD A. LOOMIS, Department of Chemistry, Washington University, One Brookings Drive, CB 1134, Saint Louis, MO 63130; ANNE B. McCOY, Department of Chemistry, The Ohio State University, Columbus, OH 43210.

Multiple features associated with transitions of the near T-shaped and linear Ne⋯I^{35,37}Cl(X,v′′=0) isomers are observed in laser-induced fluorescence spectra recorded in the ICl B−X, 2−0 and 3−0 spectral regions. The T-shaped complexes access the lowest intermolecular vibrational level within the Ne + ICl(B,v′) potential, while the linear complexes access at least four different bending levels that are delocalized in the angular coordinate. The ground state binding energy of the linear Ne⋯I^{35}Cl isomer is directly measured in two-laser, action spectroscopy experiments. A continuum signal is observed with a turn on at 17 748.4 cm^{−1} that indicates a binding energy of 84.3(2.3) cm^{−1} for the ground state complex. The ground state binding energy of the T-shaped isomer can be estimated by fitting the Ne⋯I^{35}Cl(B,v′=2) bending levels that the ground state linear isomer accesses to an ω_{e}, ω_{e}χ_{e} progression. The fit reveals a binding energy of 60.8(3.2) cm^{−1} for Ne⋯I^{35}Cl(B,v′=2), which implies a T-shaped ground state binding energy of 66.2(3.2) cm^{−1} based upon the spectroscopic shift of the T-shaped feature from the I^{35}Cl B−X, 2−0 transition. Two-laser, pump-probe experiments were also performed to access the intermolecular vibrational levels that are bound within the Ne + ICl(E 0^{+},v=10) ion-pair state. In these experiments, the pump laser was fixed on either the lowest energy level, with a T-shaped equilibrium geometry, or one of the higher energy delocalized levels within the Ne + I^{35}Cl(B,v′=2) well. In this manner, varying Ne⋯I^{35}Cl Franck-Condon windows between the B and E states are accessed and assignments of the intermolecular vibrational levels within the ion-pair state can be made. The results suggest a Ne + ICl(E 0^{+}) potential with the lowest bound level localized in the T-shaped orientation and that the higher energy levels are delocalized in the angular coordinate.